Water flows through a 7cm diameter pipe at the rate of 4meter/seconds.
1. How many cm^3 of water flows the pope in one seconds
2. Express the flow of water as a rate in liters/minute

Answer by Boreal(15235)

Cross-sectional area is pi*r^2=pi*3.5^2=12.25 pi=38.48 cm^2
4 meters/second is 400 cm/second, so the product is 38.48*400=15,392 cm^3/second
That is 15.392 liters/second or 923.52 liters/minute